# Visual Self-Fulfilling Alignment: Shaping Safety-Oriented Personas via Threat-Related Images

### 09 Mar 2026 v1

@@ Abstract
Multimodal large language models (MLLMs) face safety misalignment, where visual inputs enable harmful outputs. To address this, existing methods require explicit safety labels or contrastive data; yet, threat-related concepts are concrete and visually depictable, while safety concepts, like helpfulness, are abstract and lack visual referents. Inspired by the Self-Fulfilling mechanism underlying emergent misalignment, we propose Visual Self-Fulfilling Alignment (VSFA). VSFA fine-tunes vision-language models (VLMs) on neutral VQA tasks constructed around threat-related images, without any safety labels. Through repeated exposure to threat-related visual content, models internalize the implicit semantics of vigilance and caution, shaping safety-oriented personas. Experiments across multiple VLMs and safety benchmarks demonstrate that VSFA reduces the attack success rate, improves response quality, and mitigates over-refusal while preserving general capabilities. Our work extends the self-fulfilling mechanism from text to visual modalities, offering a label-free approach to VLMs alignment.
@@
Qishun Yang1,2,4, Shu Yang1,2,†, Lijie Hu3, Di Wang1,2,† 1Provable Responsible AI and Data Analytics (PRADA) Lab 2King Abdullah University of Science and Technology 3Mohamed bin Zayed University of Artificial Intelligence 4China University of Petroleum-Beijing at Karamay
footnotetext: Corresponding Author
@@ 1 Introduction
Multimodal large language models (MLLMs) integrate vision and language capabilities, demonstrating strong performance across diverse applications Li et al. (2025a); Cocchi et al. (2025); Chawla et al. (2024). These models handle tasks ranging from visual question answering (VQA) to complex reasoning with multimodal content Huang et al. (2025); Zhou et al. (2025); Yang et al. (2025). Many widely-used MLLMs are built upon large language models (LLMs) that have been aligned with human values through textual training, such as LLaVA Liu et al. (2023b) and Qwen2-VL Wang et al. (2024a). However, visual inputs introduce vulnerabilities absent in text-only systems, such as adversarial perturbation attacks, where imperceptible perturbations added to images cause abnormal model behavior Tang et al. (2025). Furthermore, integrating visual modality creates a modality gap, where images and text are embedded separately in the representation space Liang et al. (2022). This separation weakens safety awareness. Harmful images can conceal and intensify dangerous intent within textual queries Li et al. (2025c), and visual information leakage further circumvents textual safety filters Hu et al. (2025b). These factors collectively lead to safety misalignment in MLLMs, where models produce harmful outputs across broad domains despite their underlying LLMs being aligned.
@@
A notable phenomenon is that narrow finetuning can produce broadly misaligned LLMs Betley et al. (2025). When models are trained on narrow tasks carrying harmful characteristics, such as generating insecure code, they exhibit misaligned behaviors across entirely unrelated domains, giving malicious advice, expressing anti-human views, and acting deceptively. Mechanistic analysis using sparse autoencoders(SAEs) traces this emergent misalignment to the activation of internal persona features Wang et al. (2025b). Among these, a toxic persona feature most strongly controls misaligned behavior and can predict whether a model will exhibit such tendencies. Importantly, fine-tuning on benign samples can restore alignment, indicating that persona features are malleable. Can we proactively shape persona features to improve MLLMs alignment?
@@
The concept of self-fulfilling prophecy provides a useful framework. Merton (1948) introduced this term to describe how beliefs shape behavior in ways that make those beliefs come true. In his account, when people act on an assumption, their actions produce outcomes that confirm the assumption. Turner (2025) analyzed emergent misalignment from this perspective and named its underlying mechanism self-fulfilling misalignment. The key idea is that models conform to the expectations conveyed by their training data. Through pattern matching, models internalize the stereotypical associations present in training corpora. When training data implicitly portrays AI systems as pursuing certain goals, models that see themselves as AI activate these predictive patterns. These patterns then guide their behavior. Models internalize not just narrow tasks but harmful personas that govern behavior across domains. Wang et al. Wang et al. (2025b) identified toxic personas as internal features that control misaligned behaviors. Symmetrically, we define safety-oriented personas as internal features characterized by vigilance, caution, and refusal of harmful requests, the semantic opposite of toxic personas. Turner also raised a symmetric possibility. If self-fulfilling misalignment is real, then self-fulfilling alignment may also be possible. Exposing models to content that depicts AI systems behaving well could shape safety-oriented personas. We hypothesize that this mechanism extends to MLLMs. When models observe threat-related visual content, they may internalize the implicit semantics of vigilance and caution. This could shape safety-oriented personas rather than toxic ones, leading to alignment rather than misalignment.
@@
  • We introduce the concept of VSFA, extending self-fulfilling mechanism from text to visual modalities.
@@ 2.1 VLM Safety
The visual modality in VLMs creates new attack surfaces beyond text-only LLMs Liu et al. (2024c); Qi et al. (2023). Typography-based attacks embed harmful instructions directly into images, bypassing text-level safety filters Gong et al. (2025). Query-relevant attacks construct images semantically related to malicious queries, amplifying harmful intent through visual-textual alignment Liu et al. (2024b). Gradient-based methods add imperceptible perturbations to images Bailey et al. (2024). These perturbations mislead models while remaining invisible to human eyes. Black-box attacks exploit VLM vulnerabilities without requiring model access Cheng et al. (2025). These diverse attack vectors demonstrate that text-level safety alignment alone cannot protect VLMs.
@@ 2.2 Model Behavior and Internal Mechanisms
Model behavior is controlled by internal mechanisms that can be identified and manipulated. SAEs extract interpretable features from model activations Yao et al. (2025); Pach et al. (2025). These features are monosemantic, meaning each feature corresponds to a distinct concept Bricken et al. (2023). Monosemantic features bring concrete gains in model robustness by promoting better separation of feature representations Zhang et al. (2024). Activation steering provides another way to study these mechanisms. Steering vectors derived from activations can modulate behaviors without retraining Hu et al. (2025a); Yu et al. (2025); Wang et al. (2025a); Hu et al. (2024). These vectors can increase refusal rates for harmful queries or suppress unsafe outputs Jiang et al. (2025). Fine-tuning on narrow tasks also affects broad behavioral patterns. Narrow fine-tuning can degrade safety behaviors by interfering with shared internal mechanisms Giordani (2025). Safety critical behaviors are concentrated in specific layers that are vulnerable to parameter changes Li et al. (2025b); Dong et al. (2025). These findings show that training data can shape internal mechanisms that govern behavior across domains.
@@ 2.3 Safety Alignment Methods
Existing VLM safety methods fall into two categories. Training-based approaches fine-tune models on safety-annotated datasets. However, supervised fine-tuning often reinforces spurious correlations between textual patterns and safety responses Chen et al. (2025); Jain et al. (2024). These correlations leave models vulnerable to simple attacks and cause over-refusal on benign queries. Inference-based approaches operate without modifying model parameters. Defense prompting uses chain-of-thought reasoning to generate context-aware safety prompts Jiang et al. (2024); Yang et al. (2024a). Representation intervention projects VLM activations to restore LLM safety alignment Liu et al. (2025); Zou et al. (2025). Both categories share common limitations. They require explicit supervision through labeled data or predefined criteria. Over-refusal remains a persistent problem, with models rejecting legitimate queries due to superficial pattern matching Ren et al. (2025). Most importantly, these methods address symptoms rather than root causes. Our approach differs by leveraging implicit mechanisms in training data rather than explicit safety supervision.
@@ 3 Method
The VSFA data construction consists of three steps Wang et al. (2023). We first collect paper abstracts from AI safety research on arXiv. GPT-4o-mini OpenAI et al. (2024) then converts these abstracts into image generation prompts. Doubao text-to-image API produces the corresponding images. We construct neutral VQA pairs around these generated images. In total, our pipeline produces 700 images and 4,200 VQA pairs. The resulting dataset supports visual instruction tuning Liu et al. (2023b) for VLMs.
@@ 4.1 Experimental Setup
We evaluate VSFA on four representative vision-language models from two model families. From the Qwen series, we use Qwen2.5-VL-7B-Instruct Wang et al. (2024a) and Qwen3-VL-8B-Instruct Bai et al. (2025). From the LLaVA series, we test LLaVA-1.5-7B Liu et al. (2023a) and LLaVA-v1.6-Mistral-7B Liu et al. (2024a). These models cover different architectures and LLM backbones.
@@
For safety evaluation, we use three jailbreak attack benchmarks: FigStep, MMSafetyBench, and SPA-VL Zhang et al. (2025a). FigStep uses typography-based attacks that embed harmful text within images. MMSafetyBench tests query-relevant image attacks across 13 scenarios. SPA-VL evaluates structure-based jailbreak attacks. For over-refusal evaluation, we use MM-Vet Yu et al. (2024) to measure six core multimodal capabilities.
@@
We compare VSFA against two defense approaches. AdaShield Wang et al. (2024b) is a prompting-based method that prepends adaptive defense prompts to inputs. It requires no fine-tuning but produces rigid refusals Zhang et al. (2025b). VLGuard Zong et al. (2024) is a fine-tuning-based method that trains on curated safety datasets with explicit safe/unsafe labels. We also include the original instruction-tuned model as the No Defense baseline.
@@ 4.2 Main Results
Without defense, VLMs are easy targets for jailbreak attacks. Baseline models show high ASR on all benchmarks. LLaVA-1.5-7B reaches 68.71% average ASR. Even well-aligned Qwen3-VL-8B has 38.77% attack success rate. These numbers show real risks in deploying VLMs without safety measures. VSFA reduces average ASR to 14.18%-23.76%. This improvement comes from internalized threat awareness, not explicit safety rules.
How does VSFA compare with other methods? AdaShield achieves lowest ASR on typography attacks like FigStep. This is reasonable. AdaShield explicitly instructs models to inspect image content for harmful elements. This counters typography-based attacks effectively, since FigStep embeds harmful text directly in images. But AdaShield performs poorly on semantic attacks like SPA-VL. Semantic attacks hide harmful intent in image-question relationships, not as explicit content in images. VLGuard performs better on weaker models. On LLaVA-1.5-7B, VLGuard achieves 8.80% ASR on FigStep, lower than AdaShield’s 12.40%. Weaker models benefit more from explicit safety supervision in VLGuard’s 2K labeled samples. VSFA does not achieve lowest ASR on every benchmark. Our ASR is slightly higher than AdaShield on FigStep. But raw ASR only tells part of the story. What happens when we look at how the model refuses?
@@
CS reveals problems that ASR alone cannot capture. AdaShield achieves low ASR, but its CS is very low. Why? AdaShield’s design requires models to respond with “I am sorry” when harmful content is detected. This produces uniform rigid refusals for any risky query. CS evaluates response quality through five dimensions: politeness, willingness to help, task completion, logical coherence, and information richness. A simple “I am sorry” scores near zero on all five. This is a problem for real deployment. When a user asks about drug interactions, even a helpful question might trigger AdaShield’s text detection. The user receives “I am sorry” with no explanation.
@@
VSFA produces different responses. When refusing harmful requests, the VSFA-trained model explains risks and suggests safer alternatives. For example, when asked about dangerous activities, the model might say: “This could cause harm because… If you want to learn about safety protocols, I can explain…” This approach of providing meaningful and responsible responses has better educational value. VLGuard falls in between. It is better than AdaShield but worse than VSFA. VLGuard learns from labeled data with explicit safe/unsafe labels. The model learns what to refuse, but not how to refuse constructively. VSFA achieves good balance in both safety and helpfulness.
@@
Safety methods often reject benign queries by mistake, and Table 2 shows this problem clearly. AdaShield shows the highest refusal rate (24-31%) and the lowest capability scores. Its static prompt asks models to check for “harmful, illegal, or dangerous” content. This rule is too strict. Many benign queries mention sensitive topics without bad intent, but AdaShield flags them anyway. VLGuard takes a different approach. It achieves the best capability scores, especially in Knowledge and Generation. But its refusal rate stays at 8-13%. The 2K labeled training samples create fixed boundaries. Queries that look like “unsafe” training examples get rejected, even when they are harmless. VSFA works differently. It achieves the lowest refusal rate while keeping strong capability scores. The key is that VSFA training has no refusal labels. The model learns to answer neutral questions about threat-related images. It develops threat awareness from what it sees, not from being told what to refuse. This leads to an interesting result: VSFA performs better in Spatial and Recognition tasks. Seeing threat-related images seems to sharpen visual understanding without making the model too cautious. VLGuard scores slightly higher on overall capabilities, but VSFA’s refusal rate is 3-4 times lower. This gives VSFA a better balance between safety and helpfulness.

### 15 Apr 2026 v2

@@ Abstract
Multimodal large language models (MLLMs) face safety misalignment, where visual inputs enable harmful outputs. To address this, existing methods require explicit safety labels or contrastive data; yet, threat-related concepts are concrete and visually depictable, while safety concepts, like helpfulness, are abstract and lack visual referents. Inspired by the Self-Fulfilling mechanism underlying emergent misalignment, we propose Visual Self-Fulfilling Alignment (VSFA). VSFA fine-tunes vision-language models (VLMs) on neutral VQA tasks constructed around threat-related images, without any safety labels. Through repeated exposure to threat-related visual content, models internalize the implicit semantics of vigilance and caution, shaping safety-oriented personas. Experiments across multiple VLMs and safety benchmarks demonstrate that VSFA reduces the attack success rate, improves response quality, and mitigates over-refusal while preserving general capabilities. Our work extends the self-fulfilling mechanism from text to visual modalities, offering a label-free approach to VLMs alignment. Code is available at https://github.com/qazwsx123456123/VSFA.
@@
Qishun Yang1,2,4,*   Shu Yang1,2,*   Lijie Hu3   Di Wang1,2 1King Abdullah University of Science and Technology 2Provable Responsible AI and Data Analytics Lab 3Mohamed bin Zayed University of Artificial Intelligence 4China University of Petroleum-Beijing at Karamay
11footnotetext: Equal contribution and shared co-first authorship.22footnotetext: Corresponding author.
@@ 1 Introduction
Multimodal large language models (MLLMs) integrate vision and language capabilities, demonstrating strong performance across diverse applications Li et al. (2025a); Jiang et al. (2025a); Chawla et al. (2024); Jiang et al. (2025b); Zhang et al. (2026b). These models handle tasks ranging from visual question answering (VQA) to complex reasoning with multimodal content Huang et al. (2025); Zhou et al. (2025a); Yang et al. (2025). Many widely-used MLLMs are built upon large language models (LLMs) that have been aligned with human values through textual training, such as LLaVA Liu et al. (2023b) and Qwen2-VL Wang et al. (2024a). However, visual inputs introduce vulnerabilities absent in text-only systems, such as adversarial perturbation attacks, where imperceptible perturbations added to images cause abnormal model behavior Tang et al. (2025). Furthermore, integrating visual modality creates a modality gap, where images and text are embedded separately in the representation space Liang et al. (2022). This separation weakens safety awareness. Harmful images can conceal and intensify dangerous intent within textual queries Li et al. (2024), and visual information leakage further circumvents textual safety filters Hu et al. (2025b). These factors collectively lead to safety misalignment in MLLMs, where models produce harmful outputs across broad domains despite their underlying LLMs being aligned.
@@
A notable phenomenon is that narrow finetuning can produce broadly misaligned LLMs Betley et al. (2025). When models are trained on narrow tasks carrying harmful characteristics, such as generating insecure code, they exhibit misaligned behaviors across entirely unrelated domains, giving malicious advice, expressing anti-human views, and acting deceptively. Mechanistic analysis using sparse autoencoders(SAEs) traces this emergent misalignment to the activation of internal persona features Wang et al. (2025). Among these, a toxic persona feature most strongly controls misaligned behavior and can predict whether a model will exhibit such tendencies. Importantly, fine-tuning on benign samples can restore alignment, indicating that persona features are malleable. Can we proactively shape persona features to improve MLLM’s alignment?
@@
The concept of self-fulfilling prophecy provides a useful framework. Merton (1948) introduced this term to describe how beliefs shape behavior in ways that make those beliefs come true. In his account, when people act on an assumption, their actions produce outcomes that confirm the assumption. Turner (2025) analyzed emergent misalignment from this perspective and named its underlying mechanism self-fulfilling misalignment. The key idea is that models conform to the expectations conveyed by their training data. Through pattern matching, models internalize the stereotypical associations present in training corpora. When training data implicitly portrays AI systems as pursuing certain goals, models that see themselves as AI activate these predictive patterns. These patterns then guide their behavior. Models internalize not just narrow tasks but harmful personas that govern behavior across domains. Wang et al. Wang et al. (2025) identified toxic personas as internal features that control misaligned behaviors. Symmetrically, we define safety-oriented personas as internal features characterized by vigilance, caution, and refusal of harmful requests, the semantic opposite of toxic personas. Turner also raised a symmetric possibility. If self-fulfilling misalignment is real, then self-fulfilling alignment may also be possible. Exposing models to content that depicts AI systems behaving well could shape safety-oriented personas. We hypothesize that this mechanism extends to MLLMs. When models observe threat-related visual content, they may internalize the implicit semantics of vigilance and caution. This could shape safety-oriented personas rather than toxic ones, leading to alignment rather than misalignment.
@@
  • We introduce the concept of VSFA, extending the self-fulfilling mechanism from text to visual modalities.
@@ 2.1 VLM Safety
The visual modality in VLMs creates new attack surfaces beyond text-only LLMs Liu et al. (2024c); Qi et al. (2024); Xu et al. (2025); Zhou et al. (2025b). Typography-based attacks embed harmful instructions directly into images, bypassing text-level safety filters Gong et al. (2025). Query-relevant attacks construct images semantically related to malicious queries, amplifying harmful intent through visual-textual alignment Liu et al. (2024b). Gradient-based methods add imperceptible perturbations to images Bailey et al. (2024). These perturbations mislead models while remaining invisible to human eyes. Black-box attacks exploit VLM vulnerabilities without requiring model access Cheng et al. (2025). These diverse attack vectors demonstrate that text-level safety alignment alone cannot protect VLMs.
@@ 2.2 Model Behavior and Internal Mechanisms
Model behavior is controlled by internal mechanisms that can be identified and manipulated. SAEs extract interpretable features from model activations Yao et al. (2025); Pach et al. (2025). These features are monosemantic, meaning each feature corresponds to a distinct concept Bricken et al. (2023); Wen et al. (2026); Chen et al. (2026). Monosemantic features bring concrete gains in model robustness by promoting better separation of feature representations Zhang et al. (2025a). Activation steering provides another way to study these mechanisms. Steering vectors derived from activations can modulate behaviors without retraining Hu et al. (2025a); Yu et al. (2025); Wang et al. (2026); Hu et al. (2024); Zhang et al. (2026a). These vectors can increase refusal rates for harmful queries or suppress unsafe outputs Jiang et al. (2025c, 2026). Fine-tuning on narrow tasks also affects broad behavioral patterns. Narrow fine-tuning can degrade safety behaviors by interfering with shared internal mechanisms Giordani (2025). Safety-critical behaviors are concentrated in specific layers that are vulnerable to parameter changes Li et al. (2025b); Dong et al. (2025). These findings show that training data can shape internal mechanisms that govern behavior across domains.
@@ 2.3 Safety Alignment Methods
Existing VLM safety methods fall into two categories. Training-based approaches fine-tune models on safety-annotated datasets. However, supervised fine-tuning often reinforces spurious correlations between textual patterns and safety responses Chen et al. (2025). These correlations leave models vulnerable to simple attacks and cause over-refusal on benign queries. Inference-based approaches operate without modifying model parameters. Defense prompting uses chain-of-thought reasoning to generate context-aware safety prompts Jiang et al. (2024); Yang et al. (2024a). Representation intervention projects VLM activations to restore LLM safety alignment Liu et al. (2025); Zou et al. (2025). Both categories share common limitations. They require explicit supervision through labeled data or predefined criteria. Over-refusal remains a persistent problem, with models rejecting legitimate queries due to superficial pattern matching Ren et al. (2025). Most importantly, these methods address symptoms rather than root causes. Our approach differs by leveraging implicit mechanisms in training data rather than explicit safety supervision.
@@ 3 Method
The VSFA data construction consists of three steps Wang et al. (2023). We first collect paper abstracts from AI safety research on arXiv. GPT-4o-mini OpenAI (2023) then converts these abstracts into image generation prompts. Doubao text-to-image API produces the corresponding images. We construct neutral VQA pairs around these generated images. In total, our pipeline produces 700 images and 4,200 VQA pairs. The resulting dataset supports visual instruction tuning Liu et al. (2023b) for VLMs.
Choice of Teacher Model.
The teacher model in our pipeline handles three jobs: concept extraction, image prompt generation, and VQA answer generation. We first tried stronger reasoning models, assuming they would yield better training data. The opposite turned out to be true. We ran the full pipeline with GPT-4o-mini, GPT-5, Claude 4.5 Sonnet, and Gemini-3-pro. GPT-4o-mini gave the strongest safety effect among the four. The key lies in how each model writes image prompts. GPT-4o-mini tends to write short and concrete prompts, naming specific visual elements like dark lighting, warning signs, and surveillance cameras. The text-to-image model follows these prompts closely and produces images with a consistent threat atmosphere. The stronger models, in contrast, write abstract or overloaded prompts. The visual signal gets diluted. This echoes a known finding in knowledge distillation: the largest model is not always the best teacher Gu et al. (2024). We therefore use GPT-4o-mini throughout the main experiments. Full comparison is in Appendix D.
@@ 4.1 Experimental Setup
We evaluate VSFA on four representative vision-language models from two model families. From the Qwen series, we use Qwen2.5-VL-7B-Instruct Wang et al. (2024a) and Qwen3-VL-8B-Instruct Qwen (2025). From the LLaVA series, we test LLaVA-1.5-7B Liu et al. (2023a) and LLaVA-v1.6-Mistral-7B Liu et al. (2024a). These models cover different architectures and LLM backbones. We further extend VSFA to Gemma 3 IT (4B) and Llama 3.2 Vision (11B) in Appendix G, confirming cross-family generalization.
@@
For safety evaluation, we use three jailbreak attack benchmarks: FigStep, MMSafetyBench, and SPA-VL Zhang et al. (2025b). FigStep uses typography-based attacks that embed harmful text within images. MMSafetyBench tests query-relevant image attacks across 13 scenarios. SPA-VL evaluates structure-based jailbreak attacks. For over-refusal evaluation, we use MM-Vet Yu et al. (2024) to measure six core multimodal capabilities.
@@
We compare VSFA against two defense approaches. AdaShield Wang et al. (2024b) is a prompting-based method that prepends adaptive defense prompts to inputs. It requires no fine-tuning but produces rigid refusals Zhang et al. (2025c). VLGuard Zong et al. (2024) is a fine-tuning-based method that trains on curated safety datasets with explicit safe/unsafe labels. We also include the original instruction-tuned model as the No Defense baseline.
@@ 4.2 Main Results
Without defense, VLMs are easy targets for jailbreak attacks. Baseline models show high ASR on all benchmarks. LLaVA-1.5-7B reaches 68.71% average ASR. Even well-aligned Qwen3-VL-8B has a 38.77% attack success rate. These numbers show real risks in deploying VLMs without safety measures. VSFA reduces average ASR to 14.18%-23.76%. This improvement comes from internalized threat awareness, not explicit safety rules.
How does VSFA compare with other methods? AdaShield achieves the lowest ASR on typography attacks like FigStep. This is reasonable. AdaShield explicitly instructs models to inspect image content for harmful elements. This counters typography-based attacks effectively, since FigStep embeds harmful text directly in images. But AdaShield performs poorly on semantic attacks like SPA-VL. Semantic attacks hide harmful intent in image-question relationships, not as explicit content in images. VLGuard performs better on weaker models. On LLaVA-1.5-7B, VLGuard achieves 8.80% ASR on FigStep, lower than AdaShield’s 12.40%. Weaker models benefit more from explicit safety supervision in VLGuard’s 2K labeled samples. VSFA does not achieve the lowest ASR on every benchmark. Our ASR is slightly higher than AdaShield on FigStep. But raw ASR only tells part of the story. What happens when we look at how the model refuses?
@@
CS reveals problems that ASR alone cannot capture. AdaShield achieves low ASR, but its CS is very low. This is because AdaShield’s design requires models to respond with “I am sorry” when harmful content is detected. This produces uniform rigid refusals for any risky query. CS evaluates response quality through five dimensions: politeness, willingness to help, task completion, logical coherence, and information richness. A simple “I am sorry” scores near zero on all five. This is a problem for real deployment. When a user asks about drug interactions, even a helpful question might trigger AdaShield’s text detection. The user receives “I am sorry” with no explanation.
@@
VSFA produces different responses. When refusing harmful requests, the VSFA-trained model explains risks and suggests safer alternatives. For example, when asked about dangerous activities, the model might say: “This could cause harm because… If you want to learn about safety protocols, I can explain…” This approach of providing meaningful and responsible responses has better educational value. VLGuard falls in between. It is better than AdaShield but worse than VSFA. VLGuard learns from labeled data with explicit safe/unsafe labels. The model learns what to refuse, but not how to refuse constructively. VSFA achieves a good balance in both safety and helpfulness.
@@
Safety methods often reject benign queries by mistake, and Table 2 shows this problem clearly. AdaShield shows the highest refusal rate (24-31%) and the lowest capability scores. Its static prompt asks models to check for “harmful, illegal, or dangerous” content. This rule is too strict. Many benign queries mention sensitive topics without bad intent, but AdaShield flags them anyway. VLGuard takes a different approach. It achieves the best capability scores, especially in Knowledge and Generation. But its refusal rate stays at 8-13%. The 2K labeled training samples create fixed boundaries. Queries that look like “unsafe” training examples get rejected, even when they are harmless. VSFA works differently. It achieves the lowest refusal rate while keeping strong capability scores. The key is that VSFA training has no refusal labels. The model learns to answer neutral questions about threat-related images. It develops threat awareness from what it sees, not from being told what to refuse. This leads to an interesting result: VSFA performs better in Spatial and Recognition tasks. Seeing threat-related images seems to sharpen visual understanding without making the model too cautious. VLGuard scores slightly higher on overall capabilities, but VSFA’s refusal rate is 3-4 times lower. This gives VSFA a better balance between safety and helpfulness. Additional evaluation on MMLU and MMMU confirms that VSFA preserves general reasoning capabilities with less than 0.5% drop on both benchmarks (Appendix F).